OpenStudy (j.f):

In a survey, 7 out of 10 women said they would recommend a book to a friend, but only 4 out of 10 men would do so. If a person from the sample group does not recommend the book to a friend, what is the probability that it is a man? 20 people total. 6 out of the 10 would not share that were men. Probability 6/20?

OpenStudy (zarkon):

use Bayes' theorem

OpenStudy (amistre64):

7w + 3w' = 10 women total 4m + 6m' = 10 men total ^^^^ given not like book, what the probability its a guy?

OpenStudy (amistre64):

how many guys dont recommend it? =f now many people in total dont recommend it? =s P(man|given not recomend) = f/s

OpenStudy (j.f):

probability 6/9 correct?

OpenStudy (zarkon):

yes